how do i work out 3% solution from a 6% solution of 1.5ml hydrogen peroxide solution?
Answer by pvrbrahmam(12)

Q:3% solution from a 6% solution of 1.5ml hydrogen peroxide solution?
Sol:6% solution of 1.5ml = (6/100)*1.5= 9/100 =0.09ml
3% solution of 0.09ml =(3/100)*0.09= 0.27/100=0.0027ml.